OpenStudy (anonymous):

2x^2-8x+4=0 Solve by completing the square

OpenStudy (anonymous):

x^2-4x+2=0 x^2-4x+4-4+2=0 (x-2)^2-2=0 (x-2)^2=2 (x-2)=+,- sqrt2 x=2+,- sqrt2

OpenStudy (anonymous):

Thank you.